OpenStudy (marigirl):

e^-x cos x – e^-x sinx=0

OpenStudy (marigirl):

\[e^x(\cos(x)-\sin(x))=0\]

OpenStudy (marigirl):

how can i solve for x now Please help me :)

zepdrix (zepdrix):

The exponential is negative power? like this?\[\large\rm e^{-x}(\cos x-\sin x)=0\]

OpenStudy (marigirl):

oh yes i missed writing the negative. how can i solve it now.

zepdrix (zepdrix):

Apply your `Zero-Factor Property`: \(\large\rm (a)(b)=0\implies a=0~or~b=0\)

zepdrix (zepdrix):

So,\[\large\rm e^{-x}=0 \qquad\qquad\qquad (\cos x-\sin x)=0\]

OpenStudy (marigirl):

i am not sure how to do the cos (x)-sin(x) part

zepdrix (zepdrix):

Hmmm, I would probably add sin x to the other side. and then ummm...\[\large\rm \cos x=\sin x\]No ideas from here? :)

OpenStudy (marigirl):

cos (x)+sin(x)=1

zepdrix (zepdrix):

No no, \(\large\rm \cos^2x+\sin^2x=1\) Not \(\large\rm \cos x+\sin x=1\) :D

zepdrix (zepdrix):

Lot of ways to proceed from here. Maybe the most straight forward approach would be to divide by cosine.

zepdrix (zepdrix):

\[\large\rm 1=\frac{\sin x}{\cos x}\]

zepdrix (zepdrix):

Do you remember an identity for sine over cosine? :o

OpenStudy (marigirl):

cot(x)

OpenStudy (marigirl):

pi/4 is the answer! thanks!

zepdrix (zepdrix):

Hmm cotangent is the one with co on top :d you need tangent. but ya.. k cool :3

zepdrix (zepdrix):

You need to recall that the exponential function has an asymptote at zero. So this other part is actually never true: \(\large\rm e^{-x}=0\)

OpenStudy (marigirl):

if i divide by sin(x) ..

OpenStudy (marigirl):

oh yup